Find the solution of the Riccati equation
dy/dx=(2cos^2(x)-sin^2(x)+y^2)/2cos(x); y1(x)=sinx
1
Expert's answer
2020-03-27T12:05:03-0400

Find the solution of the Riccati equation

dy/dx=(2cos^2(x)-sin^2(x)+y^2)/2cos(x); y1(x)=sinx


Solution:

"y'=(2cos^2x-sin^2x+y^2)/2cos(x)\newline
y' =(0.5sec(x))y^2+(2cos^2(x)-sin^2(x))/2cos(x)"

"y1=sinx"


riccati eqyation general form

"y'(x) = A(x)y^2 + B(x)y + C(x)...(1)\newline
A(x) =0.5sec(x)\newline
B(x)=0\newline
C(x)=(2cos^2(x)-sin^2(x))/2cos(x)"


now substituting


"y(x)=v(x)+y_1(x)"

"y'(x)=v'(x)+y_1'(x)\newline"


in equation (1)

"y_1 \hspace{0.1cm}is \hspace{0.1cm}solution\hspace{0.1cm} to \hspace{0.1cm}above \hspace{0.1cm}differential \hspace{0.1cm}equation\hspace{0.1cm}
\newline
y

_
1' = A(x)y
^2_
1 + B(x)y_1 + C(x)."


"v'+y_1'== A(x)[v + y1]^2 + B(x)[v + y1
] + C(x)\newline
⇒ v'+ [A(x)y^2_
1 + B(x)y_1 + C(x)

] = A(x)v^2 + 2A(x)y_1v + A(x)y^
2_
1
+B(x)v + B(x)y_1 + C(x)\newline 
⇒ v' = A(x)v^
2 + 2A(x)y_1v + B(x)v\newline 
\newline
⇒ v' + [−2A(x)y_1(x) − B(x)]v = A(x)

v^2...(2)
\newline 
 \newline
 
equation\hspace{0.1cm} (2)\hspace{.1cm} is \hspace{.1cm}bernouli \hspace{.1cm} equation
\newline v'+p(x)v=q(x)v^2\newline 
p(x) = [−2A(x)y_1(x) − B(x)]=-2tan(x)
\newline q(x) = A(x)=0.5sec(x)
\newline"


solving this bernouli equation

"v'-2tan(x)v=0.5sec(x)v^2\newline
let\hspace{0.1cm}z=1/v\implies z'=-1/v^2\newline
z'+2ztan(x)=-0.5sec(x)\newline
this \hspace{0.1cm} is \hspace{0.1cm} linear \hspace{0.1cm} differential \hspace{0.1cm} equation\newline
IF=exp(\intop2tan(x)dx)=sec^2x
\newline
\therefore [sec^2(x).z]'=-0.5sec(x)\newline
z=(tan(x)sec(x)+ln|tan(x)+sec(x)|)/(-4sec^2(x))\newline
\because z=1/v\newline

 \newline
v=-4sec^2(x)/(tan(x)sec(x)+ln|tan(x)+sec(x)|)\newline
y=v+y_1\newline
y=-4sec^2(x)/(tan(x)sec(x)+ln|tan(x)+sec(x)|)+sinx"


Answer:

"y=-4sec^2(x)\/(tan(x)sec(x)+ln|tan(x)+sec(x)|)+sinx"